### v1

# Theory and Design of Space-Time Metallic Metasurfaces for Wireless Communications

Abstract
This paper details a class of metal-based space-time metasurfaces for application in wireless communications scenarios. Concretely, we describe space-time metasurfaces that periodically alternate their properties in time between three spatial states: ”air”, ”conductor” and ”grating”. We analyze the physics of these metastructures via a computationally-efficient analytical technique based on the use of Floquet-Bloch series, integral equations and circuit models. By doing so, we reveal important features of these spatiotemporal metasurfaces: scattering parameters, field profiles, diffraction angles and nature of the space-time harmonics. The results, corroborated with a self-implemented numerical FDTD approach, show the potential application of these space-time metasurfaces as beamformers acting in reflection, in transmission or both. The amplitude and direction of the diffracted orders can be electronically controlled with the paramaters of the metasurface. Moreover, the intrinsic ability of time-modulated diffractive metasurfaces to mix and multiply frequencies is tested. We show how two different modulations can lead to the same diffraction angle but with different mixed output frequencies.
Index Terms:
Space-time, time-modulation, metasurfaces, analytical method, wireless communications, beamforming, frequency mixing.
I Introduction
Metamaterials have been conceived as artificially engineered devices with the property of manipulating electromagnetic waves. The term metamaterial originally referred to structures exhibiting negative permittivity and permeability simultaneously [1]. However, in the subsequent years this conception has experimented different evolutions, remarking the emergence of thinner/planar versions called metasurfaces [2].
A conventional metasurface can be physically described as a thin screen with planar geometry exhibiting spatial modulation [3]. Thanks to this modulation, metasurfaces become excellent terminals for a wide range of applications, such as microwave absorbers [4], beam shaping [5], or lenses [6], among others. Currently, researchers are paying close attention to a new evolution of metasurfaces, the so-called space-time metasurfaces, in which time is employed as an additional modulation to the existing purely spatial one [7, 8]. This introduces additional degrees of freedom as well as richer physics on the system [9, 10, 11]. Likewise, time modulation also enhances the capabilities of the terminals, leading to promising concepts such as intelligent programmable metasurfaces [12] that could be combined and integrated into wireless communications and radars systems [13].
The recent development and fabrication of magnetic-free non-reciprocal devices [14, 15, 16, 17, 18] has been a key factor for the interest in space-time modulated devices [19, 20]. Thus, the number of theoretical studies of space-time varying structures in the microwave and optics range [21, 9, 22] have increased in the last few years favoring the emergence of novel theoretical applications, such as the use of temporal photonic crystals to achieve amplification [23], filtering and isolation [24], the utilization of travelling-wave modulations for power combining [25], the use of grounded slab for efficient phase conjugations [26], production of temporal chirping and lensing [27], or frequency mixing and multiplying [28].
Specifically, in the field of telecommunications, radar and wireless systems, frequency mixing and multiplying is of capital relevance [29, 30]. Traditionally, frequency mixing has been achieved via nonlinear circuit components such as Schottky diodes, GaAs FETs and CMOS transistors. Interestingly, frequency mixing is an intrinsic property of time-modulated diffractive systems, as the frequency of the diffracted waves is directly related to the frequency of the incident wave [31, 32]. Thus, spatiotemporal metasurfaces can complement the use of nonlinear components and even replace them in frequency ranges or scenarios where they are no longer functional. Time-modulated antennas have opened new alternatives in this regard as well [33, 34].
Furthermore, the versatility to simultaneously modify the momentum and frequency of waves has led space-time-modulated devices to be strong candidates for beamforming or beamsteering [35, 36]. Recently, it has been possible to test the beamsteering capabilities of these devices by modifying the phase constant associated to each unit cell with different temporal sequences [37]. It probes the potential of these metasurfaces as candidates for the future intelligent communications [38].
Many of these applications are usually simulated using numerical techniques such as the finite-difference time-domain (FDTD) method [39]. However, the high computational costs and long simulation times that this code entails, have prompted the development of adaptive mesh-based solutions to enhance computational efficiency, such as the Discontinuous Galerkin Time-Domain (DGTD) method [40]. Others numerical solutions based on generalized sheet transition conditions (GSTCs) [41, 42], modal techniques [43] and integral-equation methods [44] show high performance when simulating time-varying metasurfaces. Nonetheless, many times the researchers also search approaches that can provide a deep physical insight of the studied problems. Thus, some alternatives based on analytical and semi-analytical techniques have been reported in the literature [45, 46, 47, 48].
Following this line, fully-analytical circuits for 1D-spatial, 2D-spatial, 3D-spatial [49, 50, 51] and 1D-temporal [52, 53] modulations have been previously reported by the authors. Furthermore, we have recently exploited the use of Lorentz transformations to reduce a (1+1)-D spatiotemporal problem to a time-invariant (1D-spatial) system [54]. Now, in this work, we extend the circuital approach to a different set of (1+1)-D space-time scenarios. The scattering parameters, as well as information on the diffraction spectrum, are extracted in significantly shorter computation times compared to general numerical methods. In fact, the main advantage that this method presents is the possibility to simulate a wideband range of frequencies in a few seconds.
The paper is organized as follows. Section II presents the space-time metasurface and depicts the analytical framework. We discuss on the physics of the diffraction spectrum, field profiles, nature of Floquet harmonics and scattering parameters via the associated equivalent circuit. Section III focuses on describing the feasibility of the spatiotemporal metasurface for use in wireless applications. Concretely, we focus on its beamforming and frequency-mixing capabilities, both studied via the analytical method and a supportive self-implemented FDTD approach. Finally, general conclusions are drawn in Section IV.
II Analytical Framework
Figure 1: Sketch of the space-time-varying metasurface. The metasurface can alternate in time between three spatial states: air (A), conductor (C) and grating (G). By smartly combining the three states, a rich diffraction phenomenology can be created, from which wireless communications systems can take advantage.
II-A Space-Time Metasurface
Fig. 1 depicts the spatio-temporal system studied in this manuscript. The system is formed by a space-time-modulated diffractive metasurface that manipulates the incident transverse electric (TE) or transverse magnetic (TM) plane waves. The diffractive metasurface is based on different slabs painted in red and blue that alternate their electromagnetic properties in time. Each of the slabs can either become electromagnetically transparent or behave like a solid metal as it is assumed in [55].
Depending on the electromagnetic behavior of each of the slabs, the space-time-modulated diffractive metasurface can operate in three different states: “air” (A), “conductor” (C) and “grating” (G). In the air state, all slabs appears to vanish from an electromagnetic perspective, so full transmission of the incident wave is expected. In the conductor state, all slabs behave as a solid metal achieving a solid metallic wall, so full reflection of the incident wave is expected. In the case of the grating state, the red slabs behave as solid metals, and the blue slabs become electromagnetically transparent, causing the space-time metasurface to operate as a spatially-modulated metallic diffraction grating.
The geometrical parameters of the metasurface are the following. The space-time diffractive metasurface is placed along the XY plane, as Fig. 1 illustrates. In the time intervals in which the metasurface is in the grating state, it turns into a periodic arrangement of perfect electric conductor (PEC) strips (period $P$ and slit width $W$) that are infinitely-extended along the $x$ axis. The spatial periodicity of the grating state is along the $y$ axis. In the conductor and air states, the metasurface can be modeled as infinitely-extended uniform thin PEC and air screens, respectively. The air state assumes electrical parameters $\varepsilon_{r}=\mu_{r}=1$. In order to construct a proper analytical framework, we only need to enforce the temporal variation of the system to follow a time-periodic scheme, of period $T_{\mathrm{M}}$.
By wisely combining the three states in time (A-G, C-G or A-C), a rich diffraction phenomenology can be created. From an engineering perspective, the diffraction pattern can be electrically tuned by setting the space-time parameters of the system. This enables an efficient beamforming platform that could be potentially applied in wide range of wireless communications scenarios. Remarkably, the space-time metasurface can act as a beamformer in reflection, transmission or both simultaneously by simply tuning its electrical parameters. This will be illustrated in detail in the next sections.
The three states of the space-time metasurface can be realized by alternating between total transparency or total reflectivity, as recently achieved in [56]. Nonetheless, the implementation of the metasurface could also be inspired by several innovative prototypes in the microwave range [57]. Many of these devices are based on adjustable biased PIN diodes [58, 59, 60] and varactors [18, 61, 17], traditionally implemented in reflectarray and trasmittarray systems. Moreover, novel theoretical designs based on the use of electronically-reconfigurable materials such as graphene [62, 63] or transparent conductive oxides [64] open new alternatives for the millimeter-wave and terahertz regimes.
II-B Floquet-Bloch Expansion
Since the diffractive metasurface behaves as space-time-periodic media, the transverse electric and magnetic fields in regions (1) and (2) can be expressed in terms of Floquet-Bloch series. Both media are electromagnetically defined by $\varepsilon_{\text{r}}^{(i)}$ and $\mu_{\text{r}}^{(i)}$. Thus, let us consider a space-time metasurface with spatial period $P$ along the $y$-direction and temporal macroperiod $T_{\text{M}}$ [53]. It receives a plane wave impinging obliquely with angular frequency $\omega_{0}$ and TE polarization. The fields at both sides of the metasurface can be expressed in terms of a Floquet expansion of harmonics. In region (1) ($z<0$):
$\mathbf{E}_{\text{t}}^{\text{TE},(1)}(y,z,t)=\bigg{[}\text{e}^{\text{j}(\omega_{0}t-k_{0}y-\beta_{00}^{(1)}z)}+R\text{e}^{\text{j}(\omega_{0}t-k_{0}y+\beta_{00}^{(1)}z)}\\ +\displaystyle\sum_{\begin{subarray}{c}\forall mn\neq 00\end{subarray}}E_{mn}^{(1)}\text{e}^{\text{j}(\omega_{n}t-k_{m}y+\beta_{mn}^{(1)}z)}\bigg{]}\hat{\mathbf{x}}$ (1)
$\mathbf{H}_{\text{t}}^{\text{TE},(1)}(y,z,t)=\bigg{[}Y_{00}^{\text{TE},(1)}\text{e}^{\text{j}(\omega_{0}t-k_{0}y-\beta_{00}^{(1)}z)}\\ -RY_{00}^{\text{TE},(1)}\text{e}^{\text{j}(\omega_{0}t-k_{0}y+\beta_{00}^{(1)}z)}\\ -\displaystyle\sum_{\begin{subarray}{c}\forall mn\neq 00\end{subarray}}Y_{mn}^{\text{TE},(1)}E_{mn}^{(1)}\text{e}^{\text{j}(\omega_{n}t-k_{m}y+\beta_{mn}^{(1)}z)}\bigg{]}\hat{\mathbf{y}}\,.$ (2)
In the former expressions, $E_{mn}^{(1)}$ is the amplitude of the $(m,n)$-th space-time harmonic operating in region (1), $R$ is the reflection coefficient associated with the incident wave ($00$-th harmonic), $\omega_{n}$ is the angular frequency associated with the temporal $n$-th harmonic,
$\omega_{n}=\omega_{0}+n\frac{2\pi}{T_{\text{M}}},$ (3)
and $k_{m}$ is the transverse wavenumber linked to the spatial $m$-th harmonic:
$k_{m}=k_{t}+m\frac{2\pi}{P}$ (4)
with
$k_{t}=\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\,\frac{\omega_{0}}{c}\sin{(\theta_{0})}$ (5)
and $\theta_{0}$ being the incidence angle.
In a similar way, we define the tangential electromagnetic field via Floquet expansion in the region $(2)$ ($z>0$),
$\mathbf{E}_{\text{t}}^{\text{TE},(2)}(y,z,t)=\bigg{[}T\text{e}^{\text{j}(\omega_{0}t-k_{0}y-\beta_{00}^{(2)}z)}\\ +\displaystyle\sum_{\begin{subarray}{c}\forall mn\neq 00\end{subarray}}E_{mn}^{(2)}\text{e}^{\text{j}(\omega_{n}t-k_{m}y-\beta_{mn}^{(2)}z)}\bigg{]}\hat{\mathbf{x}}$ (6)
$\mathbf{H}_{\text{t}}^{\text{TE},(2)}(y,z,t)=\bigg{[}TY_{00}^{\text{TE},(2)}\text{e}^{\text{j}(\omega_{0}t-k_{0}y-\beta_{00}^{(2)}z)}\\ +\displaystyle\sum_{\begin{subarray}{c}\forall mn\neq 00\end{subarray}}Y_{mn}^{\text{TE},(2)}E_{mn}^{(2)}\text{e}^{\text{j}(\omega_{n}t-k_{m}y-\beta_{mn}^{(2)}z)}\bigg{]}\hat{\mathbf{y}}$ (7)
where $T$ denotes the transmission coefficient and $E_{mn}^{(2)}$ the amplitude associated to the $(m,n)$-th Floquet harmonic. In the magnetic fields expansions in (2) and (7), $Y_{mn}^{(i)}$ denote the admittance of the the $(m,n)$-th Floquet harmonic located at the $i$-th medium, defined for TE incidence as
$Y_{mn}^{\text{TE},(i)}=\frac{\beta_{mn}^{(i)}}{\mu_{r}^{(i)}\mu_{0}\,\omega_{n}},\hskip 14.22636pt$ (8)
with
$\beta_{mn}^{(i)}$ $=\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}\left[\frac{\omega_{n}}{c}\right]^{2}-\left[k_{m}\right]^{2}}\,.$ (9)
being the propagation constant of the $(m,n)$-th harmonic in region $(i)$.
The previous derivation can easily adapted for TM incidence. In the TM case, the electric and magnetic field orientations are the opposite to the TE case, thus
$\mathbf{E}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{x}}\cdot\mathbf{E}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\cdot\hat{\mathbf{y}}$ (10)
$\mathbf{H}_{\text{t}}^{\text{TM},(i)}(y,z,t)$ $=[\hat{\mathbf{y}}\cdot\mathbf{H}_{\text{t}}^{\text{TE},(i)}(y,z,t)]\cdot\hat{\mathbf{x}}\,.$ (11)
In addition, the admittance expressions must now be replaced to those related to TM harmonics,
$Y_{mn}^{\text{TM},(i)}=\frac{\varepsilon_{r}^{(i)}\varepsilon_{0}\,\omega_{n}}{\beta_{mn}^{(i)}}\,.$ (12)
II-C Nature of the Space-Time Harmonics
Real(imaginary) values of $\beta_{mn}^{(i)}$ imply that the nature of the $(m,n)$-th Floquet harmonic is propagative(evanescent). By using eq. (9), it can be shown that all evanescent space-time harmonics fulfill the general conditions
$\frac{\sqrt{\varepsilon_{r}\mu_{r}}}{c}\left(\omega_{0}\sin(\theta_{0})-\left|\omega_{0}+n\frac{2\pi}{T_{\text{M}}}\right|\right)+m\frac{2\pi}{P}>0$ (13a)
$\frac{\sqrt{\varepsilon_{r}\mu_{r}}}{c}\left(\omega_{0}\sin(\theta_{0})+\left|\omega_{0}+n\frac{2\pi}{T_{\text{M}}}\right|\right)+m\frac{2\pi}{P}<0$ (13b)
In a simpler scenario where the surrounding media are air ($\varepsilon_{r}=\mu_{r}=1$) and the waves impinge normally ($\theta_{0}=0$) to the space-time metasurface, eq. (II-C) reduces to:
$-\left|1+n\frac{T_{0}}{T_{\text{M}}}\right|+m\frac{\lambda_{0}}{P}>0$ (14a)
$+\left|1+n\frac{T_{0}}{T_{\text{M}}}\right|+m\frac{\lambda_{0}}{P}<0$ (14b)
Figure 2: Propagative (black) and evanescent (white) nature of $(m,n)$-harmonics when an incident wave of angular frequency $\omega_{0}$ = $2\pi\cdot 30$ GHz impinges on the space-time-varying metasurface with: (a) $P$ = $0.7\lambda_{0}$ and $T_{\text{M}}$ = $2T_{0}$, (b) $P$ = $0.3\lambda_{0}$ and $T_{\text{M}}$ = $2T_{0}$, (c) $P$ = $0.7\lambda_{0}$ and $T_{\text{M}}$ = $3T_{0}$. Normal incidence and $\varepsilon_{r}=\mu_{r}=1$ are assumed.
The former expressions show that, for a fixed temporal macroperiod $T_{\text{M}}$, there will be a greater number of evanescent space-time harmonics as the spatial periodicity $P$ reduces. Likewise, for a fixed $P$, there will be a greater number of evanescent harmonics as $T_{\text{M}}$ increases. Moreover, all the $(0,n)$-th harmonics ($m=0,\forall n$) are propagative under normal incidence. This is consistent with the results extracted in our previous work [52], as the $(0,n)$-th Floquet harmonics are only driven by the time modulation.
Fig. 2 summarizes all the aforementioned phenomenology related to the space-time harmonics. In the figure, the vertical axis represents the spatial ($m$-indexed) harmonics and the horizontal axis represents the temporal ($n$-indexed) harmonics. The evanescent harmonics are marked in white and the propagative ones in black. Upon analyzing Fig. 2, it is evident the existence of certain symmetry for both indexes. In the case of considering normal incidence, this symmetry appears around the null of the propagation constant in both the vertical and horizontal planes. Therefore, seeking the roots of eq. (9) under normal incidence, the $(m,n)$-th harmonic with $\beta_{mn}=0$ emerges for $m=0$ and $n=-\frac{T_{\text{M}}}{T_{0}}$. It corresponds to a DC harmonic ($\omega_{n}=0$) [42], which is present at ($m=0,n=-2$) in Fig. 2(a)-(b) and at ($m=0,n=-3$) in Fig. 2(c).
The phenomenology of the propagative and evanescent harmonics is much richer here than in the case considered in our previous works [52, 53]. In this work, we are considering a space-time modulation, where in [52, 53] we considered a time-only modulation. Time-only modulations led to scenarios where most of the harmonics are propagative, and only a few are evanescent. In fact, it was shown in [52] that all time harmonics are propagative under normal incidence conditions in the time-modulated metasurface. The situation is rather different and more interesting from a physics and engineering perspective when spatial and temporal modulations intervene together [65].
II-D Diffraction Angles and Beamforming
The diffraction angle of the $(m,n)$-th Floquet harmonic can be computed using:
$\theta_{mn}^{(i)}=\mathrm{asin}\left(\dfrac{k_{m}}{\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}}\left[\frac{\omega_{n}}{c}\right]}\right)=\mathrm{atan}\left(\frac{k_{m}}{\beta_{mn}}\right)$ (15a)
$\theta_{mn}^{(i)}=\mathrm{asin}\left(\dfrac{\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\text{sin}(\theta_{0})+m\frac{\lambda_{0}}{P}}{\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}}\left[1+n\frac{T_{0}}{T_{\text{M}}}\right]}\right)$ (15b)
At a first sight, one can visualize in eq. (15a) that propagation below the XZ plane, i.e., negative angles $\theta_{mn}^{(i)}$, is allowed. A negative transverse wavenumber $k_{m}$ (with real-valued $\beta_{mn}^{(i)}$) is easily achievable for some negative indexes $m$, which translates into a negative diffraction angle $\theta_{mn}^{(i)}$. This constitutes a major difference with respect to our previous works on time-only metasurfaces, where negatives diffraction angles were forbidden [52]. Additionally, eq. (15b) reveals that most of the diffraction harmonics tend to close to the normal ($\theta_{mn}\rightarrow 0$) when the temporal period $T_{\text{M}}$ decreases or, analogously, when the time-modulation frequency $\omega_{\text{M}}$ increases, with the exception of the $(m,0)$-th harmonics that are not affected by the time modulation. Therefore, the diffraction pattern produced here is of a much richer nature. Thus, the beamforming capabilities of the present space-time metasurface are expected to significantly exceed those of the previous space-only and time-only configurations.
II-E Reflection/Transmission Coefficients
The derivation of the scattering parameters, i.e., the reflection and transmission coefficients, require of prior knowledge of the electric-field profiles at the space-time discontinuity. This is an insightful technique that has been successfully applied in the past for the analysis of 1D-spatial [66, 49], 1D-time [52, 53], 2D-spatial [67, 68, 69] and 3D-spatial [70] periodic systems. Formally, the space-time profile, also referred to as base function (bf), can be described by the function $\mathbf{E}_{\mathrm{bf}}(y,t)$ depending on space and time.
The application of the continuity of the electric fields through the space-time interface ($z=0$),
$\mathbf{E}_{t}^{\text{TE/TM},(1)}(y,0,t)=\mathbf{E}_{t}^{\text{TE/TM},(2)}(y,0,t)=\mathbf{E}_{\mathrm{bf}}(y,t)\,,$ (16)
leads to
$1+R$ $=T$ (17)
$1+R$ $=\frac{\widetilde{E}_{\mathrm{bf}}(k_{0},\omega_{0})}{PT_{\text{M}}}$ (18)
$E_{mn}^{(1)}$ $=E_{mn}^{(2)}=\frac{\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})}{PT_{\text{M}}}\,,$ (19)
with
$\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})=\displaystyle\int_{-\frac{P}{2}}^{\frac{P}{2}}\displaystyle\int_{0}^{T_{\text{M}}}E_{\mathrm{bf}}(y,t)\,\text{e}^{-\text{j}(\omega_{n}t+k_{m}y)}\text{d}y\,\text{d}t\,.$ (20)
Note that ${E}_{\mathrm{bf}}(y,t)$ is simply the instantaneous amplitude of the vector $\mathbf{E}_{\mathrm{bf}}(y,t)$ representing the base function and $\widetilde{E}_{\mathrm{bf}}(\omega_{n},k_{m})$ is its Fourier transform.
From eqs. (18) and (19), we may also obtain that
$E_{mn}^{(1)}=E_{mn}^{(2)}=(1+R)N(k_{m},\omega_{n})\,,$ (21)
where
$N(k_{m},\omega_{n})=\frac{\widetilde{E}_{\mathrm{bf}}(k_{m},\omega_{n})}{\widetilde{E}_{\mathrm{bf}}(k_{0},\omega_{0})}$ (22)
is a term that accounts for the coupling between the fundamental $00$-th harmonic and the corresponding $(m,n)$-th space-time harmonic.
Now, the continuity of the instantaneous Poynting vector is imposed at the space-time interface ($z=0$). The power passing through the interface is evaluated over a temporal period $T_{\text{M}}$ and over a spatial period $P$. This leads to
$\int_{-\frac{P}{2}}^{\frac{P}{2}}\displaystyle\int_{0}^{T_{\text{M}}}\mathbf{E}_{\mathrm{bf}}(y,t)\times\mathbf{H_{t}}^{\text{TE/TM},(1)}(y,0,t)\,\text{d}y\,\text{d}t\\ =\displaystyle\int_{-\frac{P}{2}}^{\frac{P}{2}}\displaystyle\int_{0}^{T_{\text{M}}}\mathbf{E}_{\mathrm{bf}}(y,t)\times\mathbf{H_{t}}^{\text{TE/TM},(2)}(y,0,t)\,\text{d}y\,\text{d}t\,$ (23)
By replacing the values of eqs. (2), (7) (TE incidence) or eq. (11) (TM incidence) and eq. (21) into eq. (23) and operating, we achieve:
$(1-R)Y_{00}^{\text{TE/TM},(1)}-(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(1)}\,|N(k_{m},\omega_{n})|^{2}\\ =(1+R)Y_{00}^{\text{TE/TM},(2)}+(1+R)\displaystyle\sum_{\begin{subarray}{c}mn\neq 00\end{subarray}}Y_{mn}^{\text{TE/TM},(2)}|N(k_{m},\omega_{n})|^{2}\,,$ (24)
By manipulating (24), the reflection coefficient $R$ admits to be expressed as
$R=\frac{Y_{00}^{\text{TE/TM},(1)}-Y_{00}^{\text{TE/TM},(2)}-Y_{\mathrm{eq}}}{Y_{00}^{\text{TE/TM},(1)}+Y_{00}^{\text{TE/TM},(2)}+Y_{\mathrm{eq}}}\,.$ (25)
The term $Y_{\mathrm{eq}}$, which groups the double sums in eq. (24), is the equivalent admittance that models the space-time metasurface. It is computed as
$Y_{\mathrm{eq}}=\displaystyle\sum_{\begin{subarray}{c}\forall mn\neq 00\end{subarray}}\left|N(k_{m},\omega_{n})\right|^{2}\left(Y_{mn}^{\text{TE/TM},(1)}+Y_{mn}^{\text{TE/TM},(2)}\right).$ (26)
II-F Circuit Topology
Figure 3: Circuit topology that describes the physics of the space-time-varying metasurface: (a) Complete equivalent circuit including the two-dimensional sum of infinite transmission lines representing each $(m,n)$-th harmonic. (b) Reduced equivalent circuit for TE and TM incidences.
Eqs. (25) and (26) are circuitally interpreted by the topology shown in Fig. 3. In general, eq. (25) reveals that the circuit model associated to the analytical framework consists of two semi-infinite transmission lines modeling the input (1) and output (2) media, and an equivalent admittance $Y_{\mathrm{eq}}$ that models the space-time metasurface [see Fig. 3(b)]. A deeper insight into eq. (26) shows that the equivalent admittance is internally formed by a sum of the admittances (parallel connections of semi-infinite transmission lines) associated to each space-time Floquet harmonic.
Each of the parameters $N(k_{m},\omega_{n})$ taking part in eq. (26) is interpreted as a complex transformer that takes into consideration the coupling between the $(m,n)$-th harmonic and the fundamental ($00$-th) one. Each admittance of each Floquet harmonic is loaded with a complex transformer, except for the fundamental harmonic whose value is $N(k_{0},\omega_{0})=1$. This is sketched in Fig. 3(a). Note that, for the sake of visualization, only a few harmonics of the double infinite sum are represented in Fig. 3(a), and that the transmission lines of the fundamental harmonic are marked in blue.
At a first sight, the circuit topologies shown in Figs. 3(a) and 3(b) coincide with those shown in previous works on space-only and time-only metamaterial configurations [71, 52]. However, the value of the forming admittances and complex transformers, which rule the diffraction behavior of the metasurface, are completely different. In (three-dimensional) space-only systems, the complex transformers are a function of the transverse wavenumbers $k_{m},k_{n}$ and $k_{l}$. In time-only problems, those are a function of the angular frequency $\omega_{n}$. In spatiotemporal systems, the complex transformers are a function of the transverse wavenumbers plus the angular frequency.
It was shown that higher-order harmonics in space-only and time-only metasurfaces described by the present topology are of evanescent and propagative nature, respectively. This causes that higher-order harmonics contribute with capacitive and/or inductive terms in space-only (patch-based and aperture-based) metasurfaces [71], while higher-order harmonics contribute with a purely resistive term in time-only configurations [52]. The scenario is rather more complex in the present space-time metasurface. Higher-order harmonics fulfilling $|n|\gg|m|$ are predominantly temporal. Their propagation constant is frequency-independent and real-valued. It can be approximated as
$\beta_{mn}^{(i)}\left|{}_{|n|\gg|m|}\right.\approx\left|\frac{2\pi n}{cT_{\text{M}}}\right|\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}}\,.$ (27)
By using eqs. (12) and (8), the admittances of the $(m,n)$-th harmonic read in this case:
$Y_{mn}^{\text{TM/TE},(i)}\left|{}_{|n|\gg|m|}\right.\approx Y_{00}\,\sqrt{\frac{\varepsilon_{r}^{(i)}}{\mu_{r}^{(i)}}}$ (28)
Therefore, all the space-time harmonics with $|n|\gg|m|$ contribute together with a pure resistive term.
On the other hand, higher-order harmonics fulfilling $|m|\gg|n|$ are predominantly spatial. Their propagation constant is frequency-independent and imaginary:
$\beta_{mn}^{(i)}\left|{}_{|m|\gg|n|}\right.\approx-\text{j}\frac{2\pi}{P}\left|m\right|\,.$ (29)
The associated admittances read:
$Y_{mn}^{\text{TM},(i)}\left|{}_{|m|\gg|n|}\right.\approx\text{j}(\omega_{0}+n\omega_{\text{M}})\frac{\varepsilon_{r}^{(i)}\varepsilon_{0}P}{2\pi|m|}$ (30a)
$Y_{mn}^{\text{TE},(i)}\left|{}_{|m|\gg|n|}\right.\approx\frac{1}{j(\omega_{0}+n\omega_{\text{M}})}\,\frac{2\pi|m|}{\mu_{r}^{(i)}\mu_{0}P}$ (30b)
where $\omega_{\text{M}}=2\pi/T_{\text{M}}$ is simply the angular frequency of the time modulation.
From the former expressions, it can be identified that the $(m,n)$-th higher-order TM harmonic fulfilling $|m|\gg|n|$ contributes with two parallel capacitors. The first capacitor is dependent on the frequency of the incident wave $\omega_{0}$ and second one on the time modulation of the space-time metasurface $\omega_{\text{M}}$. When considering higher-order terms ($|n|\gg 1$), the latter capacitor is the dominant, being the time modulation of the metasurface truly relevant in this case. Similarly, the $(m,n)$-th higher-order TE harmonic fulfilling $|m|\gg|n|$ contributes with two series inductors, the one associated to the frequency $\omega_{\text{M}}$ being the dominant. Therefore, space-time harmonics with $|m|\gg|n|$ contribute jointly with a pure capacitive or inductive term, depending on the type of wave incidence.
The propagation constant of higher-order harmonics that fulfill $|m|\sim|n|$ ($|m|,|n|\gg 1$) is still frequency-independent, but cannot lead to a simple classification of the harmonic contributions into purely resistive or purely capacitive/inductive. Depending on the surrounding media ($\varepsilon_{r}$ and $\mu_{r}$) and electrical parameters of the space-time modulation ($P$ and $T_{\text{M}}$), the propagation constant $\beta_{mn}^{(i)}$ will be real or imaginary and the $(m,n)$-th higher-order harmonic fulfilling $|m|\sim|n|$ will either contribute with a resistive or a capacitive/inductive term.
Conversely, low-order harmonics have associated a propagation constant that is frequency-dependent and a function of the incident angle $\theta_{0}$, among other parameters. As a consequence, low-order terms cannot be decoupled into purely spatial or purely temporal. Thus, they cannot be directly associated to resistors, capacitors or inductors. To maintain the wideband behavior of the analytical approach, low-order harmonics should be analyzed directly with their complex admittances $Y_{mn}^{\text{TE/TM},(i)}$. Therefore, the whole equivalent admittance $Y_{\text{eq}}$ in eq. (26) can be divided into three main parallel contributions: a complex and frequency-dependent admittance $Y_{\text{eq}}^{\text{(lo)}}$ describing the effect of low-order harmonics, a resistor $R^{\text{(hi)}}$ modeling predominantly-temporal higher-order TM/TE harmonics, and either a capacitor $C^{\text{(hi)}}$ (TM incidence) or an inductor $L^{\text{(hi)}}$ (TE incidence) modeling predominantly-spatial higher-order harmonics.
If needed, further approximations can be made in the quasi-static regime (low frequencies: $\omega_{0}\ll 2\pi/T_{\text{M}}$ and $\omega_{0}/c\ll 2\pi/P$) in order to neutralize the frequency-dependence of the complex low-order admittance $Y_{\text{eq}}^{\text{lo}}$. However, this comes at the expense of reducing the operation bandwidth of the wideband analytical model. In the quasi-static regime, the time modulation truly marks the behavior of the space-time harmonics, i.e., $\omega_{n}\approx n\omega_{\text{M}}$ and $k_{m}\approx 2\pi m/P$. Thus, the admittance of the $(m,n)$-th harmonic under a quasi-static (qs) approximation is
$Y_{mn}^{\text{TM},(i),qs}\approx\frac{\varepsilon_{r}^{(i)}\varepsilon_{0}\,n\,\omega_{\text{M}}}{\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}(n\,\omega_{\text{M}})^{2}-\left(\frac{2\pi m}{P}\right)^{2}}}$ (31a)
$Y_{mn}^{\text{TE},(i),qs}\approx\frac{\sqrt{\varepsilon_{r}^{(i)}\mu_{r}^{(i)}(n\,\omega_{\text{M}})^{2}-\left(\frac{2\pi m}{P}\right)^{2}}}{\mu_{r}^{(i)}\mu_{0}\,n\,\omega_{\text{M}}}\,,$ (31b)
which can only be purely real or reactive depending on the selected $(m,n)$-th harmonic. As a consequence, the equivalent admittance in eq. (26) can be reduced in the quasi-static case to a parallel connection of a resistor $R^{\text{(hi), qs}}$ and a capacitor $C^{\text{(hi), qs}}$ (TM incidence) or the resistor $R^{\text{(hi), qs}}$ and an inductor $L^{\text{(hi), qs}}$ (TE incidence).
II-G Basis Functions
Figure 4: (a) Sketch of the C-G metasurface and the $\mathbf{E}_{\mathrm{bf}}(y,t)$ on a single space-time unit cell of the C-G configuration under TE/TM normal incidence. (b) Sketch of the A-G metasurface and the $\mathbf{E}_{\mathrm{bf}}(y,t)$ on a single space-time unit cell of the A-G configuration under TE/TM normal incidence. Parameters: $T_{\text{M}}=4T_{\text{0}}$, $t_{C}=t_{A}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $P=0.7\lambda_{0}$ and $W=0.5P$.
As detailed in Section II-E, the knowledge of the tangential electric field profile $\mathbf{E}_{\mathrm{bf}}(y,t)$ is crucial for the analysis of the space-time metasurface via the analytical method. In this subsection, we describe the analytical expressions for the space-time base functions involving the three states under consideration: air (A), conductor (C) and grating (G).
II-G1 Air state (A)
In the air state, the space-time metasurface vanishes from an electromagnetic perspective. Therefore, the base function that models the air state coincides with the electric profile of the incident plane wave; namely,
$E_{\mathrm{bf}}(y,t)=A\sin(\omega_{0}t),\,\,\forall y\in\bigg{[}-\frac{P}{2},\frac{P}{2}\bigg{]},t\in t_{A},$ (32)
where $t_{A}$ is the time in which the space-time metasurface is in the air state.
II-G2 Conductor state (C)
In the conductor state, the space-time metasurface turns into an infinitesimally-thin metallic sheet. If losses are neglected, the metallic sheet can be modeled with PEC condictions, so the tangential electric field must be zero at its surface. This leads to the base function
$E_{\mathrm{bf}}(y,t)=0,\,\,\forall y\in\bigg{[}-\frac{P}{2},\frac{P}{2}\bigg{]},t\in t_{C},$ (33)
where $t_{C}$ is the time in which the metasurface is in the conductor state.
II-G3 Grating (G)
The grating state can be modeled with the following base function
$E_{\text{bf}}(y,t)=A\sin(\omega_{0}t)\begin{cases}\Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{\frac{1}{2}},\quad&\text{TE polarization}\\ \Bigg{[}1-\bigg{(}\frac{2y}{W}\bigg{)}^{2}\Bigg{]}^{-\frac{1}{2}},\quad&\text{TM polarization}\,,\\ \end{cases}$ (34)
defined in the slit region ($-W/2\leq y\leq W/2$) and the time interval $t\in t_{G}$, time in which the space-time metasurface is in the grating state. This base function relates the field excitation and charge distribution in a conventional one-dimensional diffraction grating illuminated by a sinusoidal plane wave of frequency $\omega_{0}$.
As previously introduced, a wise combination (A-G, C-G) of the three states in a time-periodic scheme can lead to a rich diffraction phenomenology from which beamformers can benefit. Top panel of Fig. 4(a) sketches the scenario where the space-time metasurface alternates between conductor and grating states (C-G case). As shown, the scheme is periodic in space and time, with periods $P$ and $T_{\mathrm{M}}$, respectively. Moreover, the terms $t_{C}$ and $t_{G}$ indicate the time in which the metasurface is in the conductor and grating states. Note that $T_{\mathrm{M}}=t_{C}+t_{G}$ in this case to retain time periodicity. The basis function that models the C-G scenario is plotted at the bottom of Fig. 4(a) for TE and TM incident polarizations. Null regions indicate that the space-time metasurface is in the conductor state [eq. (33)], while $\cap$- and $\cup$-shaped regions [eq. (34)] are associated to the grating state under TE and TM incidences, respectively.
Similar rationale is applied to the air-conductor (A-G) case, sketched at left side of Fig. 4(b). The A-G configuration is periodic in space and time, with periods $P$ and $T_{\mathrm{M}}$, respectively. In this case, the time period is constituted by two addends, $T_{\mathrm{M}}=t_{A}+t_{G}$, related to the time in which the metasurface is in the air and grating states. The basis function of the A-G case is illustrated in the right side of Fig. 4(b). $\cap$/$\cup$-shaped regions indicate that the space-time metasurface is in grating state, while sinusoidal regions indicate that the metasurface is in the air state [eq. (32)].
Transients between air, conductor and grating states are neglected along the document. In general, the electric field at the space-time metasurface takes a nonzero time in reaching the steady state for the considered metasurface state. However, transient phenomena has no appreciable impact on the diffraction pattern as long as the considered operating frequencies are sufficiently low. In our particular case, we must ensure operation far below the plasma frequency $\omega_{p}$ of the metal under consideration and time periods ($T_{0}$ and $T_{\text{M}}$) much greater than the relaxation time $\tau$ of the free charges. Metals that are good conductors such as copper, aluminum, silver or gold have plasma frequencies of the order of $\omega_{p}\sim 10^{15}$ rad/s and relaxation times $\tau\sim 10^{-14}$ s. Thus, transient times are several orders of magnitude shorter than the periods of the incident wave and time modulation in the radio, microwave and millimeter-wave frequencies. In practice, the operation frequency of the analytical approach is limited down to the low-THz regime.
The different states of the structure can also be invoked by regarding reconfigurable metasurfaces, which for this case would consist of a reconfigurable grating. Reconfiguration is realized via active elements such as diodes or varactors [72], or via electronic materials such as graphene [69] or $\text{Ge}_{2}\text{Sb}_{2}\text{Te}_{2}$ [73]. Their electronic effect over the grating is tuned to vary periodically, invoking periodic cycles in which the states of full transparency (A), full reflectivity (C), or the natural grating (G) interchange in the sense described above. This can be possible thanks to the excitation of resonant states in the grating through these elements. The existence of a transient regime when passing from one state to the other is neglected since we are assuming transient time to be temporarily short enough in comparison with the rest of times ($t_{C},t_{A},t_{G}$) taking part in a periodic cycle. Small transient times have negligible effects on the Fourier Transform of the electric field profile at the discontinuity [18].
The considered basis functions also limit the frequency range of the analytical approach. Specifically, the most constraining basis function is the one that models the grating state. It was reported in previous works on spatially-periodic metasurfaces that the maximum operating frequency is limited, from a practical perspective, to $f_{\text{max}}\sim 1.5c/P$ [74, 75]. Thus, the proposed analytical works in a wide frequency range, even within the grating lobe regime. In the case of considering oblique incidence, the analytical approach is expected to work fine as well for both TM and TE polarization states, but the maximum operating frequency could reduce to $f_{\text{max}}\sim c/P$ for large incident angles. Moreover, the slit width should be less than $W\lesssim 0.7P$ to obtain accurate results. That is, the analytical approach is wideband, robust to large incidence angles and not strictly constrained to the subwavelength regime. Taking into account the aforementioned limitations, from an engineering perspective, we will be able to excite many propagative harmonics if we go into the diffraction regime with $P\gg w$. This scenario would lead to a richer diffraction pattern.
III Wireless Applications
In this section, we test the capabilities of the space-time metasurface for wireless applications. We show how the electrical tuning of the different space-time parameters can lead to a rich diffraction phenomenology, mixing frequencies, setting specific angular regions and favoring transmission or reflection. Furthermore, we validate the present analytical approach by comparing the results with full-wave simulations in a self-implemented finite-difference time-domain (FDTD) code.
III-A Harmonic Distribution
We start by considering the case in which the space-time metasurface alternates between the conductor and grating states. The basis function that models the C-G case is mathematically described by a combination of eqs. (33) and (34). Fig. 5 depicts the obtained normalized amplitudes of the Floquet harmonics under normal TE and TM incidences. In all these simulations, the incidence of a monochromatic wave with a frequency of $30$ GHz has been assumed. The results are obtained with the proposed analytical model and then compared with our FDTD code. The agreement between the analytical and FDTD results is good, thus validating the present approach.
The harmonic pattern shows similarities for TE and TM incidences. Note that two main lobes appear in Fig. 5 for $n=0$ and $n=-16$. These indexes correspond to the frequencies $\omega_{n}=\pm\omega_{0}$, respectively. However, in the TM case, the amplitude of the spatial ($m$-indexed) harmonics show a slower decay; namely, a larger number of spatial modes are needed to accurately reconstruct the basis function. This is due to the singularity of the spatial profile present in the TM case that is not in the TE case. On the other hand, the harmonic pattern under TE incidence is mainly governed by the temporal $n$-indexed harmonics.
Fig. 5(b) presents the Floquet harmonic pattern in the case of reducing the time period from $T_{\text{M}}=8T_{0}$ to $T_{\text{M}}=2T_{0}$. The peaks are observed for the temporal harmonics $n=0$ and $n=-4$. As illustrated, the two main lobes approach each other in the case of reducing the time period (increasing the switching frequency of the metasurface). The spatial and temporal harmonic dependencies are decoupled, as eqs. (32)-(34) impose and Fig. 5 confirms. Thus, the behaviour of the spatial modes for both incidences is not affected by modifying the time period.
Figure 5: Normalized amplitude of the Floquet coefficients when varying the temporal period in a C-G scenario. Normal TM and TE incidences are considered. Cases: (a) $T_{\text{M}}=8T_{0}$ and (b) $T_{\text{M}}=2T_{0}$. The basis function $E_{\mathrm{bf}}(y,t)$ is plotted along the results. Parameters: $t_{C}=[0,\ 0.5T_{\text{M}}]$, $t_{G}=[0.5T_{\text{M}},\ T_{\text{M}}]$, $P=0.7\lambda_{0}$, $W=0.5P$.
Figure 6: Normalized amplitude of the Floquet coefficients in a C-G scenario. Normal TM incidence is assumed and the basis function $E_{\mathrm{bf}}(y,t)$ is plotted along the results. (a) The duty cycle is varied. First row: $D=0.25$. Second row: $D=0.75$. Here, $D$ and $1-D$ indicates the amount of time that the metasurface is in the conductor and grating states, respectively. Parameters: $T_{\text{M}}=8T_{0}$, $P=0.7\lambda_{0}$ and $W=0.5P$. (b) The spatial ratio $W/P$ is varied. First row: $W=0.1P$. Second row: $W=0.5P$. Parameters: $T_{\text{M}}=4T_{0}$, $P=0.7\lambda_{0}$ and $D=0.5$.
Another temporal parameter of interest is the duty cycle $D$ [53]. The duty cycle is a dimensionless parameter that relates the amount of time that the space-time metasurface is in each of the three states. As it will be discussed later, the duty cycle plays an important role in controlling the amplitude of the diffracted, reflected and transmitted, waves. For the present C-G case, the metasurface keeps in conductor state a time given by $DT_{\mathrm{M}}$, with $T_{\text{M}}$ being the temporal macroperiod. Otherwise, it remains in the grating state $(1-D)T_{\mathrm{M}}$ seconds. First and second row of Fig. 6(a) show the harmonic pattern when the duty cycle is $D=0.25$ and $D=0.75$, respectively. Increasing the duty case in this case implies a greater presence of the grating state. Therefore, the width of the main lobes narrows along the temporal ($n$-indexed) axis as the spatiotemporal field profile (basis function) is more similar to that of a pure sine.
The last spatial parameter of interest is the relation between the slit width $W$ and the spatial period $P$; namely, the dimensionless ratio $W/P$. Logically, decreasing the slit width in a grating implies an increase in the reflection through the structure. From a harmonic perspective [see the harmonic pattern in Fig. 6(b)], decreasing the ratio $W/P$ provokes that the width of the main lobes widens. This is specially noticeable in the TM case, as the basis function includes the presence of two singularities at the edge of the slits (positions $y=\pm W/2$ within the unit cell) in a $\cup$-shaped profile. The two singularities get closer as $W/P$ reduces. As a consequence, more spatial modes are needed to reconstruct the tangential electric-field profile due to the rapid spatial variation of the basis function along $y$.
III-B Diffracted Waves
Figure 7: FDTD simulation of the diffracted electric fields for the (a) A-G and the (b) C-G configurations. The space-time metasurface is placed at $z/\lambda_{0}=20$. Oblique TE and TM incidences ($\theta_{0}=40^{\text{o}}$) are analyzed. Parameters: $T_{\text{M}}=4T_{0}$, $t_{C}=t_{A}=[0,\ 0.5T_{\text{M}}]$, $t_{G}=[0.5T_{\text{M}},\ T_{\text{M}}]$, $P=0.7\lambda_{0}$, $W=0.5P$.
To gain a more comprehensive understanding on the diffraction patterns of these metasurfaces, Fig. 7(a) and Fig. 7(b) illustrate the electric-field distributions of the A-G and C-G cases, respectively, extracted with the self-implemented FDTD code. The same duty cycle $D=0.5$ is maintained in both configurations for a fair comparison.
In general, higher transmission is observed for TM polarization than for TE polarization. This is in line with the results obtained in previous works on spatially-modulated surfaces [66, 49]. However, when the metasurface alternates between air and grating states [A-G configuration in Fig. 7(a)], a high degree of transmission is generally seen. In contrast, when the metasurface switches between the conductor and grating states [C-G configuration in Fig. 7(b)], high reflection is created. This is a somewhat expected but interesting feature from an engineering perspective, since the air and conductor states tend to favor transmission and reflection, respectively. This feature of the space-time metasurface, combined with the tuning of the duty cycle parameter, can be exploited in practice for the efficient design of reflection-based, transmission-based or mixed reflection-transmission-based beamformers.
From a physical perspective, it is noteworthy that the incorporation of spatial modulation by the grating leads to the emergence of spatial modes in the upper ($y\geq 0$) and lower ($y<0$) half-spaces. This constitutes a major difference with respect to our previous works on time-only modulated metamaterials, where only diffraction in the upper half-space was permitted [52, 53]. Thus, the present space-time metasurface creates a richer diffraction pattern compared to time-only configurations.
III-C Frequency Mixing
Figure 8: FDTD results for a C-G frequency mixer for an incident wave of frequency $\omega_{0}=30$ GHz: (a) $T_{\text{M}}=T_{0}$; (b) $T^{\prime}_{\text{M}}=2T_{0}$; (c) $T^{\prime}_{\text{M}}=3T_{0}$. Normal TM incidence is considered. Parameters: $t_{C}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $P=1.2\lambda_{0}$, $W=0.5P$.
Frequency mixing is vital for some physics and engineering applications such as radio astronomy, telecommunications or radar systems. Traditionally, nonlinear components such as Schottky diodes, GaAs FETs and CMOS transistors are used for this purpose. In this regard, temporal and spatiotemporal metamaterials possess the inherent ability to mix frequencies [32]. This is essentially due to “temporal” diffraction and the production of higher-order Floquet harmonics ruled by the condition $\omega_{n}=\omega_{0}+2\pi n/T_{\text{M}}$ [see eq. (3)]. Thus, the proposed space-time metasurface can produce new frequencies, $\omega_{n}$, from the frequency of the incident wave, $\omega_{0}$.
Interestingly, eq. (15a) suggests that two different space-time modulations can lead to the same diffraction angle, but with different output frequencies, by simply tuning their space-time parameters. In order to show this frequency-mixing effect, let us consider the two following modulations. Modulation #1 is defined by the (unprimed) space-time parameters $\{m,n,P,T_{\text{M}}\}$. Similarly, modulation #2 is defined by the (primed) parameters $\{m^{\prime},n^{\prime},P^{\prime},T^{\prime}_{\text{M}}\}$. This leads to the diffraction angles $\theta_{mn}$ and $\theta_{m^{\prime}n^{\prime}}$ (associated to modulations #1 and #2, respectively), given by
$\theta_{mn}=\text{asin}\left(\frac{k_{t}+\frac{2\pi m}{P}}{\frac{2\pi}{cT_{0}}\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\Big{(}1+n\frac{T_{0}}{T_{\text{M}}}\Big{)}}\right)$ (35)
$\theta_{m^{\prime}n^{\prime}}=\text{asin}\left(\frac{k_{t}+\frac{2\pi m^{\prime}}{P^{\prime}}}{\frac{2\pi}{cT_{0}}\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\Big{(}1+n^{\prime}\frac{T_{0}}{T^{\prime}_{\text{M}}}\Big{)}}\right)$ (36)
Nonetheless, both configurations are expected to have the same spatial period ($P^{\prime}=P$), since this parameter cannot be electronically reconfigured in practice like the duty cycle $D$ or the time period $T_{\text{M}}$. Under this assumption, by equating eqs. (35) and (36), we have
$\frac{k_{t}+\frac{2\pi m}{P}}{\frac{2\pi}{cT_{0}}\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\Big{(}1+n\frac{T_{0}}{T_{\text{M}}}\Big{)}}=\frac{k_{t}+\frac{2\pi m^{\prime}}{P}}{\frac{2\pi}{cT_{0}}\sqrt{\varepsilon_{r}^{(1)}\mu_{r}^{(1)}}\Big{(}1+n^{\prime}\frac{T_{0}}{T_{\text{M}}^{\prime}}\Big{)}}$ (37)
The former equation reduces under normal incidence ($k_{t}=0$) to
$\frac{m}{m^{\prime}}=\frac{1+n\frac{T_{0}}{T_{\text{M}}}}{1+n^{\prime}\frac{T_{0}}{T_{\text{M}}^{\prime}}}$ (38)
Expressions (37) and (38) give us the design equations to produce frequency mixing between harmonics of different orders ($\{m,n\}$ and $\{m^{\prime},n^{\prime}\}$) while maintaining the same output diffraction angle. This can be done by simply adjusting the temporal periods of the two modulations, $T_{\text{M}}$ and $T^{\prime}_{\text{M}}$.
Fig. 8 shows a practical example of frequency mixing. For simplicity, we have selected an scenario ruled by normally-incident waves. However, the previous discussion can be straightforwardly extended to the design of spatiotemporal frequency mixers that operate under oblique-incidence conditions by using eq. (37). In the figure, a monochromatic plane wave of frequency $\omega_{0}$ impinges normally on a C-G space-time metasurface (located at $z/\lambda_{0}=0$). As seen, the metasurface creates a rich diffraction pattern, where only the transmitted diffracted fields are plotted for a better visualization of the results. With the design information given by eq. (38), the temporal periods of Figs. 8(a), 8(b), 8(c) have been carefully chosen. The underlying idea is to keep the same diffraction angle while varying the output frequencies in the three configurations that are shown. For instance, it can be seen that the diffraction orders ($m=2,n=2$) and ($m^{\prime}=1,n^{\prime}=1$) in Figs. 8(a) and 8(b) share the same diffraction angle ($\theta_{22}=\theta_{1^{\prime}1^{\prime}}=33.72^{\mathrm{o}}$) while their output frequencies are different ($\omega_{2}=3\omega_{0}$, $\omega_{1^{\prime}}=1.5\omega_{0}$). The aforementioned diffraction orders are marked with a blue line in Figs. 8(a) and 8(b). Similarly, the diffraction orders ($m=3,n=3$) and ($m^{\prime}=1,n^{\prime}=1$), marked with a black line in Figs. 8(a) and 8(c), have the same diffraction angle ($\theta_{33}=\theta_{1^{\prime}1^{\prime}}=38.65^{\mathrm{o}}$) while their output frequencies are $\omega_{3}=4\omega_{0}$ and $\omega_{1^{\prime}}=1.33\omega_{0}$, respectively. As this example seeks to illustrate, the capabilities of the present metal-based metasurface are promising for the production of frequency mixers and frequency multipliers based on space-time architectures that could be integrated in wireless communication systems.
III-D Transmissive Beamformer
Figure 9: (a)-(b) Diffraction patterns of the A-G configuration under oblique TM incidence ($\theta_{0}=40^{\mathrm{o}}$): (a) $T_{\text{M}}=4T_{\text{0}}$ and $P=0.7\lambda_{0}$. (b) Different spatial period: $P=\lambda_{0}$. Fixed parameters: $t_{A}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $W=0.5P$. (c)-(d) Diffraction patterns of the A-G configuration under normal TE incidence: (c) $T_{\text{M}}=4T_{\text{0}}$ and $P=1.2\lambda_{0}$. (d) Different temporal period: $T_{\text{M}}=T_{0}$. Fixed parameters: $t_{A}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $W=0.6P$.
The main purpose of this section is to exploit the performance of these structures as transmission beamformers. Thus, we test the A-G configuration by changing the spatial and temporal periods. For a better visualization in the simulations, only the transmissive part of the diagrams is plotted. The A-G metasurface is placed at the beginning of the space simulation ($z/\lambda_{0}=0$). The most important diffraction orders have been marked with an arrow to facilitate their visualization. and their space and time indexes $(m,n)$ have been included to know their nature.
Thus, Fig. 9(a) depicts the diffraction pattern of a monochromatic wave that impinges in the spatiotemporal metasurface under oblique TM incidence ($\theta_{0}=40^{\text{o}}$). Note how, in the upper part of the transmission plane ($y/\lambda_{0}>0$), the diffraction angles must be similar to the time-only case since the $(m=0,n)$-th harmonics are implied. However, in this manuscript, the beamformer capabilities have been increased since the arising of the new space-time harmonics in the lower part of the transmission diagram ($y/\lambda_{0}<0$). In these examples, it is noticeable the emerging of the $(-1,n)$-th harmonics.
Fig. 9(b) shows the same scenario but with a slight increase of spatial period (from $P=0.7\lambda_{0}$ to $P=\lambda_{0}$). The rest of the parameters, $T_{\text{M}}=4T_{0}$, $W=0.5P$ and $D=0.25$, are kept fixed. Fig. 9(b) illustrates how $(m=-1,n)$-th harmonics approach the normal as the spatial period $P$ and the temporal index $n$ increase. This is well predicted by eq. (15b). Moreover, notice how the temporal $(m=0,n)$-th harmonics are not affected by a change in the spatial period, since the temporal period is the same in both Figs. 9(a) and 9(b).
In contrast, Figs. 9(c)-(d) depict the phenomenology when the temporal period is modified. For simplicity, TE normal incidence is assumed and slit width is increased to $W=0.6P$ in order to achieve a higher transmission. Furthermore, a spatial period greater than the wavelength of the incident wave is imposed to enrich the diffraction diagram since the first (spatial) grating lobes are excited. Fig. 9(c) shows the diffraction pattern for the same time period as considered throughout the previous section of the manuscript ($T_{\text{M}}=4T_{0}$). In this case, a symmetric pattern is appreciated where the temporal $(0,n)$-th harmonics are in the same direction ($\theta_{0n}=0^{\mathrm{o}}$). Moreover, all the $(0,n)$-th harmonics are propagative due to their temporal nature and the normal incidence conditions. However, since spatial modulation has been included, new spatiotemporal harmonics emerge symmetrically at both the upper and lower parts of the diagram. Finally, in Fig. 9(d), it is observed how the higher-order $(m,n)$-th harmonics approach to the normal when the temporal period is reduced (from $T_{\text{M}}=4T_{0}$ to $T_{\text{M}}=T_{0}$), as it was predicted in Section II-D.
Table I provides a comparison of the diffraction angles for the previous cases shown in Fig. 9. They have been extracted with the analytical approach [by means of eq. (15b)] and the FDTD method. A good agreement is appreciated between the analytical and numerical results.
TABLE I: Diffraction angles obtained for each $(m,n)$-th harmonic using the FDTD and analytical method for the cases shown in Fig. 9.
Incidence ${\theta_{mn}}({{}^{\mathrm{o}}})$ Analytical FDTD
${T_{\text{M}}=}4T_{0}$ ${P=}0.7\lambda_{0}$ $\theta_{0-1}$ $58.98$ $58.17$
$\theta_{00}$ $40.00$ $40.36$
$\theta_{01}$ $30.94$ $31.27$
$\theta_{02}$ $25.37$ $25.40$
${\theta_{0}^{\text{TM}}=}40^{\mathrm{o}}$ $\theta_{03}$ $21.54$ $20.38$
$\theta_{-13}$ $-26.64$ $-25.98$
$\theta_{-12}$ $-31.54$ $-30.20$
$\theta_{-1-1}$ $-38.89$ $-39.43$
$\theta_{-10}$ $-51.70$ $-52.40$
${T_{\text{M}}=}4T_{0}$ ${P=}\lambda_{0}$ $\theta_{0-1}$ $58.98$ $58.17$
$\theta_{00}$ $40.00$ $40.36$
$\theta_{01}$ $30.94$ $31.27$
$\theta_{02}$ $25.37$ $25.40$
${\theta_{0}^{\text{TM}}=}40^{\mathrm{o}}$ $\theta_{03}$ $21.54$ $20.38$
$\theta_{-11}$ $-16.57$ $-15.28$
$\theta_{-10}$ $-20.88$ $-21.18$
$\theta_{-1-1}$ $-28.38$ $-31.38$
$\theta_{-1-2}$ $-45.48$ $-41.50$
${T_{\text{M}}=}4T_{0}$ ${P=}1.2\lambda_{0}$ $\theta_{\pm 10}$$\pm 56.38$$\pm 56.37$
${\theta_{0}^{\text{TE}}=}0^{\mathrm{o}}$ $\theta_{\pm 11}$$\pm 41.77$$\pm 42.61$
$\theta_{\pm 12}$$\pm 33.72$$\pm 34.41$
$\theta_{\pm 13}$$\pm 28.41$ $28.7$
${T_{\text{M}}=}T_{0}$ ${P=}1.2\lambda_{0}$ $\theta_{\pm 10}$$\pm 56.38$$\pm 55.98$
$\theta_{\pm 44}$$\pm 41.77$$\pm 43.15$
${\theta_{0}^{\text{TE}}=}0^{\mathrm{o}}$ $\theta_{\pm 22}$$\pm 33.72$$\pm 34.31$
$\theta_{\pm 46}$$\pm 28.41$ $27.99$
$\theta_{\pm 23}$$\pm 24.60$ $23.82$
Frequency-mixing phenomena can be appreciated in Figs. 9(c) and 9(d) as well. As discussed in Section III-C, two different space-time modulations can lead to different output frequencies while maintaining the same diffraction angle. This is case for the diffraction orders ($m=1$, $n=1$) in Fig. 9(c) and ($m^{\prime}=4$, $n^{\prime}=4$) in Fig. 9(d). Both diffraction orders have associated the same diffraction angle ($\theta_{11}=\theta_{4^{\prime}4^{\prime}}=41.77^{\mathrm{o}}$) but different output frequencies ($\omega_{n=1}=1.25\omega_{0}$, $\omega_{n^{\prime}=4}=5\omega_{0}$). In a general case, eqs. (37) and (38) can be applied in the design of frequency mixers under oblique and normal incidence, respectively, by simply tuning the space-time parameters of the modulation. The key parameter in design is the time period, as the spatial period cannot be easily reconfigured electronically in the current metastructure. The present discussion opens the way to the design of spatiotemporal frequency mixers based on aperiodic metallic structures, which are expected to provide more degrees of freedom in design and ease of control than purely-periodic metallic ones [32].
III-E Reflective Beamformer
Figure 10: Reflection coefficient as a function of the frequency of the incident wave for the C-G configuration. (a) When varying the duty cycle. (b) When varying the spatial dimension ratio. TE oblique incidence $(\theta_{0}=20^{\mathrm{o}})$ is considered. Fixed parameters: $P=10$ mm, $T_{\text{M}}=4T_{0}$.
In this subsection we test the tunability of the reflection coefficient for the C-G beamformer. Fig. 10 depicts the analysis of the reflection coefficient as a function of the frequency of the incident wave when the temporal $T_{\text{M}}=4T_{0}$ ($f_{0}=4f_{\text{M}}$) and spatial $P=10$ mm periods are fixed. Note that, in the present scenario, frequency of the modulation increases as the frequency of the incident wave does. Thus, the ratio $f_{0}/f_{\text{M}}$ is fixed to a value of four.
The main advantage that the analytical approach contributes is the faster extraction of the scattering parameter for different points in frequency. Note that, the simulations carried out by means of the FDTD method [assuming a monochromatic impinging wave] takes several minutes depending on the mesh and the simulation space. However, the circuital approach takes only a few seconds to simulate the whole frequency range.
Fig. 10(a) shows the curves obtained by the proposed analytical method for the reflection coefficient $|S_{11}|$ when varying the duty cycle $D$. In this case, $DT_{\text{M}}$ represents the time in which the C-G metasurface is in the conductor state. FDTD simulations of the electric-field distributions are also included as insets. The black, blue and red solid lines are associated to the duty cycle $D=0.75$, $D=0.5$ and $D=0.25$, respectively. It can be seen how the amplitude of the reflection coefficient decreases as the space-time metasurface remains less time in the conductor (fully reflective) state; namely, as $D$ decreases. This is in line with the former theoretical discussion and with the FDTD results inserted in Fig. 10(a). Moreover, the oblique incident angle ($\theta_{0}=20^{\mathrm{o}}$) leads to the emerging of the grating lobe approximately at 23 GHz. This phenomenology is well caught by the circuital approach. It can be also appreciated in the FDTD results, as the grating lobes make appearance at 30 GHz (see the lower half space) but do not show at 20 GHz, below the grating-lobe regime. Finally, Fig. 10(b) shows the amplitude of the reflection coefficient in the case of varying the spatial ratio $W/P$ and fixing the duty cycle. In this case, less reflection is observed as the spatial ratio (slit width $W$) increases. As Figs. 10(a) and (10)(b) illustrate,
A crucial difference related to these two parameters should be noted. The modification of $D$ and $W$ is useful to adjust the reflection coefficient of the system, as explained in this section. However, as discussed in Section III-A, the harmonic distribution of the diffraction pattern is also affected. Nevertheless, the majority of $(m,n)$-th harmonics that can attain significant amplitudes when varying $W$ are of evanescent nature. This is not an issue from the beamforming perspective, since this parameter remains fixed in the grating state. For this reason, this parameter is the best choice for setting the desired reflection coefficient of the system. On the other hand, when $D$ is modified, many of these $(m,n)$th harmonics that can attain significant amplitudes will be propagative. Thus, this property of the temporal parameters turns them into the best option for adjusting the beamsteering of the system.
IV Conclusions
In this manuscript, we have introduced a novel metal-based spatiotemporal metasurface for application in wireless communications systems. The metasurface can alternate its properties in time within three different states: ”air” (fully transparent), ”conductor” (fully reflective) and ”grating” (partially transparent and reflective). By combining the three states in a time-periodic scheme, a rich diffraction pattern is created, which can be electronically reconfigured by setting the space-time parameters of the system.
The physics of the space-time metasurface is described by means of an analytical technique based on equivalent circuits, Floquet-Bloch expansions and integral equations. We have considered scenarios where the metasurface is illuminated by oblique TE and TM plane waves. The analytical technique has proven to be computationally efficient compared to general numerical approaches and the self-implemented FDTD code. Moreover, it provides physical insight on the diffraction spectrum, nature (evanescent/propagative) of the space-time Floquet harmonics, and scattering parameters.
The results of the work show that the present space-time metasurface offer clear advantages compared to our previous time-only configuration. Here, we have exploited the inherent ability of the space-time metasurface to mix and multiply frequencies. We have shown that two or more output frequencies can be engineered so their diffraction angles is the same. Finally, we have shown that efficient beamforming can be realized by tuning the space-time parameters of the system. Combinations of the conductor and grating states are prone to smartly reflect most of the diffracted power. Conversely, combinations of the air and grating states can enable effiecient beamforming in transmission.
Acknowledgments
This work has been supported by grant PID2020-112545RB-C54 funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It has also been supported by grants PDC2022-133900-I00, TED2021-129938B-I00 and TED2021-131699B-I00, and by Ministerio de Universidades and the European Union NextGenerationEU, under Programa Margarita Salas, by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR, and by IJC2020-043599-I funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEUPRTR.